# Learning Task Representations from In-Context Learning

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable proficiency in in-context learning (ICL), where models adapt to new tasks through example-based prompts without requiring parameter updates. However, understanding how tasks are internally encoded and generalized remains a challenge. To address some of the empirical and technical gaps in the literature, we introduce an automated formulation for encoding task information in ICL prompts as a function of attention heads within the transformer architecture. This approach computes a single task vector as a weighted sum of attention heads, with the weights optimized causally via gradient descent. Our findings show that existing methods fail to generalize effectively to modalities beyond text. In response, we also design a benchmark to evaluate whether a task vector can preserve task fidelity in functional regression tasks. The proposed method successfully extracts task-specific information from in-context demonstrations and excels in both text and regression tasks, demonstrating its generalizability across modalities. Moreover, ablation studies show that our method’s effectiveness stems from aligning the distribution of the last hidden state with that of an optimally performing in-context-learned model.
@@ 1 Introduction
Large language models (LLMs) based on the transformer architecture [50] have seen dramatic improvements in recent years. A notable feature of these models, such as GPT-3 [3], is their capability for in-context learning (ICL). This process involves the model receiving a prompt that includes demonstrations of a task in the form of input-output pairs. When presented with a new query input, the model can generate the appropriate output by extrapolating from the provided examples. For instance, after being prompted with a few examples, these models are capable of producing the antonyms of given input words. A concrete example is
$\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{vanish}\rightarrow\text{appear},\text{ short}\rightarrow\text{tall},}}_{\text{examples}}\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{increase}\rightarrow}}_{\text{query}}\underbrace{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{decrease}}}_{\text{completion}},$ (1)
@@
where the blue text is the prompt and the red text is the completion provided by the model. Not limited to linguistic tasks, transformers can also in-context learn a general class of functions $\mathcal{F}$ [8]. Specifically, for any function $f\in\mathcal{F}$, the model is capable of approximating $f(x_{\text{query}})$ for a new query input $x_{\text{query}}$. This class may include linear or nonlinear relationships, which can be represented by various machine learning models such as linear regression, multi-layer ReLU networks, and decision trees.
@@
The ICL capability is particularly intriguing as it allows models to adapt to a wide range of downstream tasks on-the-fly—i.e., without requiring any parameter updates post-training [3]. This indicates that models can extract abstract task information (the relationship between input-output pairs) from the prompt and use it to generate accurate responses for query inputs. However, due to the complex nature of the transformer architecture, the computational mechanisms that facilitate how models internally extract and encode task information are not well understood.
@@
Task vectors have been identified as key mechanisms for encoding task information during ICL [48, 25, 11, 29]. These vectors capture how models process prompts and embed the abstract in-context task as a numerical representation. While the effectiveness of these vectors is empirically well-supported, the precise computation of a task vector within the transformer architecture remains debated. Early findings suggest that layer or attention activations significantly influence ICL performance [11, 29, 48]. However, no consensus exists on the optimal way to conceptualize or compute these vectors, resulting in varied methodologies often restricted to a single modality (e.g., only text, not numeric functions). Notably, existing approaches often fail to generalize to functional regression tasks, consisting of a range of linear and nonlinear functions. Furthermore, they require adaptation to individual transformer architectures. For instance, the layers over which the task vector is computed must be explicitly identified for each model [25, 48, 11].
@@
These limitations raise concerns about the generalizability and practical utility of existing task vector formulations. Consequently, in the pursuit of a more standardized and automated pipeline for task encodings, we hypothesize that:
@@ A New Benchmark for Regression Tasks
While prior work has explored where in-context tasks are represented, most studies focus solely on language, with limited attention to functional regression. To evaluate task encodings in regression tasks, we introduce a new benchmark motivated by recent findings on the role of positional information in reasoning tasks [61]. This setup offers an intuitive framework for analyzing task encodings and ICL capabilities in regression.
@@
Our method performs effectively on the proposed regression benchmark and demonstrates significant advantages in language. Ablation studies indicate that this performance results from aligning the distribution of the last hidden state with that of an optimally in-context-learned model, providing insights into how task steering is achieved.
@@ 2 Related Work
Initial studies on developing task representations for transformers are documented in [23, 41, 33, 36, 17]. These works proposed methods to create compositional task encodings through model perturbations in the parameter space: “soft” prompts, codebooks, and meta-mappings. Notably, the term task vectors was first introduced in [17]. In contrast, through the application of causal mediation analysis [37, 51, 30, 57, 9], function vectors were discovered to exist inherently within the transformer architecture and to exhibit strong causal effects [48]. This finding aligns with research on RNNs, where it was demonstrated that RNN hidden states can be grouped based on task similarities [22, 13].
Efforts to represent ICL tasks often derive a task vector from the layer activations associated with the dummy token following the input query [11]. This approach was later refined by averaging the layer activations of dummy tokens across a few-shot prompt and optimizing them [25]. However, these methods were primarily designed for language tasks, where input-output pairs are explicitly separated by specific tokens (e.g., $\rightarrow$”). This limitation restricts their applicability to broader domains, such as regression tasks. Another approach leverages the first principal component of the difference in layer activations to guide the ICL task, resulting in In-Context Vectors (ICVs) [29]. Notably, attention heads have been argued to play a critical role in transferring information between token positions [50, 7]. Building on this insight, the Function Vector (FV) [48] is computed as the sum of activations from a specific subset of attention heads, selected based on an indirect metric derived from causal inference literature [37]. In this study, we aim to develop a structured and automated method for extracting function vectors from transformer architectures, expanding on the approach of Todd et al. [48].
@@ 3 Technical Preliminaries
$\operatorname{MultiHead}_{\ell}(Q_{\ell},K_{\ell},V_{\ell})=\operatorname{Concat}(\operatorname{head}_{\ell,1},\ldots,\operatorname{head}_{\ell,J})W_{\ell}^{O},$
$\text{where}\quad\mathbb{R}^{q}\ni\operatorname{head}_{\ell,j}\coloneqq\operatorname{Attn}_{\ell,j}=\operatorname{softmax}\big{(}Q_{\ell,j}K^{\top}_{\ell,j}/\sqrt{d_{k}})V_{\ell,j}.$
Here $W_{\ell}^{O}\in\mathbb{R}^{Jq\times d}$ is the output projection matrix, and $Q_{\ell,j}$, $K_{\ell,j}$, and $V_{\ell,j}$ are the query, key, and value matrices for each attention head $j\leq J$ at layer $\ell$. The term $\sqrt{d_{k}}$ normalizes the softmax operation for stability, where $d_{k}$ is the dimension of the key matrix. This multi-head approach allows the model to dynamically adjust its focus across different parts of the input based on the context.
Note that each head at layer $\ell$ operates in a low-dimensional space $q<d$, distinct from the main hidden state residual stream of the transformer. As observed in [6], the original transformer formulation [50] can be seen as dividing the matrix into a block form $[W_{\ell,1},\ldots,W_{\ell,J}]$, followed by directly projecting each $\operatorname{head}{\ell,j}$ into the residual space. Consequently, the attention head output $a{\ell,j}$ can be defined using the notations from [50] and [6] as:
@@ 3.1 In-Context Learning
Regression tasks refer to a function class $\mathcal{F}$ that is learned through ICL. We closely follow the formulation proposed in [8] for training and testing models on regression data. For each prompt, a random function $f$ from $\mathcal{F}$ is sampled according to a distribution $\mathcal{D}_{\mathcal{F}}$, and a set of random inputs $x_{i}\in\mathbb{R}^{m}$ for $i=1,\ldots,T$ is drawn from $\mathcal{D}_{\mathcal{X}}$. These inputs are then evaluated by $f$ to produce the prompt $p^{f}=\{x_{1},f(x_{1}),\ldots,x_{T},f(x_{T}),x_{\text{query}}\}$. We denote the model’s output as $M_{\theta}(p^{f})$.
A function class $\mathcal{F}$ can be either linear or nonlinear. For linear functions, inputs are drawn from an isotropic Gaussian distribution $\mathcal{N}(0,I_{m})$, and a random function is defined by sampling a weight vector $w$ from $\mathcal{N}(0,I_{m})$, setting $f(x)=w^{\top}x$. Here, the task is characterized by the weight vector $w\in\mathbb{R}^{m}$. For nonlinear functions, possible forms of $f$ include multi-layer ReLU networks or decision trees. We employ the models pre-trained by Garg et al. [8], with the training procedure and additional details provided in Appendix C.2.2.
@@
In language tasks, we focus on straightforward natural language processing (NLP) applications such as antonym generation, where an example is shown in (1). The model $M_{\theta}$ processes each prompts $p^{t}$ and produces a next-token distribution $M_{\theta}(\cdot\mid p^{t})$ over the words in vocabulary $\mathcal{V}$. During inference, we evaluate pre-trained LMs rather than training or fine-tuning them. Following [48], we generate corrupted prompts $\tilde{p}_{i}^{t}$, where the labels within each prompt are shuffled. This shuffling breaks the direct link between the inputs $x_{i,k}$ and the outputs $\tilde{y}_{i,k}$, leaving the provided examples uninformative about the task and effectively “blocking” ICL.
Transformer       Transformer + FV       Transformer + ICV
(a) Linear functions
(b) Sparse linear functions
(c) 2-layer ReLU NN
@@ 4.1 Motivational Observations
Our study begins by observing that existing ICL task formulations are either inapplicable [25, 11] (see Appendix C.2.6 for details) or fail to generalize to regression data [29, 48]. To demonstrate the latter, we deliberately prevent the LLM from extracting task understanding from in-context examples. If similar ICL performance is observed when the task vector is integrated into the model, it indicates that the task vector encodes the task information. In the language domain, this approach involves shuffling the labels of in-context example pairs [48], as described in Section 3.1. However, this method is unsuitable for regression tasks due to fundamental differences in how input-output relationships are represented and learned. We discuss this in detail in Appendix B.
We draw on a recent finding that highlights the critical role of positional information in reasoning tasks within ICL [61]. Since regression tasks require the model to infer and apply implicit mathematical relationships from the given examples, they can be considered instances of numeric reasoning. “Blocking” ICL in regression tasks can thus be achieved by manipulating (absolute) positional encodings through controlled sequence lengths during training and evaluation. The model is initialized and pre-trained to accommodate sequences up to a maximum length $T_{\text{max}}$, ensuring it can process longer sequences. However, during pre-training, the model is exposed only to sequences with a maximum length $T_{\text{train}}<T_{\text{max}}$, leaving positional encodings for positions beyond $T_{\text{train}}$ untrained.
At test time, we evaluate the model on sequences longer than $T_{\text{train}}$ but within $T_{\text{max}}$. This approach isolates the effects of untrained positional encodings, disrupting the model’s reasoning ability and effectively blocking ICL [61, Theorem 3]. Consequently, we expect that if a task representation effectively encodes the task, it should maintain task fidelity beyond $T_{\text{train}}$. The results, shown in Figure 1, reveal that the tested approaches fail to generalize across the evaluated class of functions. While these methods perform well in the language domain, their shortcomings on regression tasks raise concerns about the effectiveness and generalizability of their approach.
As our investigation deepens, we articulate the central research questions:
@@
    How can we develop a task representation in a principled manner that is generalizable across different modalities, such as language or mathematical functions?
@@
    How can this new representation be leveraged to improve performance on downstream tasks?
@@ 4.2 Learnable Task Vector
We introduce a causal framework for learning task-specific representations from in-context data, independent of modality and model architecture. Our approach addresses key limitations of previous methods, such as Function Vectors, by automatically computing task formulations without the need for prior model analysis or manual adjustments.
@@
The FV formulation assumes equal contributions from all attention heads to task representation by summing their activations. However, we argue that the influence of each head varies—some may contribute significantly to task encoding, while others have a lesser impact. Constraining weights to unity is therefore impractical, and focusing exclusively on a subset of attention heads risks overlooking subtle contributions from others. To address this, attention heads should be weighted in the summation to accurately capture their varying contributions.
@@
Previous work typically integrates task encodings into models by adding vectors to the outputs of specific hidden layers through simple vector summation [48, 11, 25]. However, this approach assumes uniform applicability across layers, ignoring that transformers’ hidden states represent progressively refined representations of the input. To overcome this limitation, we compute task vectors uniquely for each hidden layer, aligning them with the evolving representations. While initializing a unique set of attention head weights for each of the $L$ layers is an option, it would be computationally expensive. Instead, we compute layer-specific task vectors as a weighted sum of the attention heads within each layer:
@@
4.3 How to Learn Task Representations
We optimize LTV weights through gradient steps to approximate the “true” weights the model uses to represent ICL tasks, aiming to validate this formulation as a proof of concept. Consequently, this approach is not primarily intended to address issues like length generalization shown in Figure 1. An effective strategy for learning the parameters $\Phi$ involves optimizing the LTV in scenarios where the model (with frozen parameters) fails to perform ICL, isolating the causal contribution of $\Phi$ to the task performance. This approach ensures our method is both data-driven and causal, as it directly attributes improvements to $|\Phi|=L\times J$ parameters, facilitating application to downstream tasks. The overall pipeline for training an LTV is summarized in Algorithm 1.
Regression Tasks
$\underset{\Phi}{\text{min}}\text{ }\mathbb{E}_{f\sim\mathcal{D}_{\mathcal{F}},x\sim\mathcal{D}_{\mathcal{X}}}\bigg{[}\Big{(}\tilde{M}_{\theta}\big{(}p^{f}=\{x_{1},f(x_{1}),\ldots,x_{T_{\mathbf{v}}},f(x_{T_{\mathbf{v}}}),x_{\text{query}}\}\big{.}\mid\mathbf{v}_{\Phi}^{f}\big{)}-f(x_{\text{query}})\Big{)}^{2}\bigg{]},$ (3)
where $\mathbf{v}_{\Phi}^{f}=\{v_{\ell}^{f}\}_{\ell=1}^{L}$ represents the set of layer-wise LTVs computed for function $f$, and $\tilde{M}_{\theta}(\cdot\mid\mathbf{v}_{\Phi}^{f})$ denotes the transformer output modified by integrating $\mathbf{v}_{\Phi}^{f}$ into the corresponding hidden layers.
Language Tasks
minΦ $\underset{\Phi}{\text{min}}\text{ }\mathbb{E}_{\tilde{p}^{t}\sim\tilde{P}^{t}}\Big{[}-\log\Big{(}\tilde{M}_{\theta}\big{(}y_{\text{query}}\mid\tilde{p}^{t};\mathbf{v}_{\Phi}^{t}\big{)}\Big{)}\Big{]},$ (4)
@@ 4.3 How to Learn Task Representations
1:Input: Model $M_{\theta}$, task $t$, and number of samples $N$
@@
3:Collect: $N$ prompts $\{p^{t}_{i}\}_{i=1}^{N}$ demonstrating task $t$
@@
7:         Forward Pass: $\hat{y}_{\text{query},i}\leftarrow M_{\theta}(p^{t}_{i}\mid\mathbf{v}_{\Phi}^{t})$
8:         Compute Loss: $\mathcal{L}(y_{\text{query},i},\hat{y}_{\text{query},i})$ based on (3) or (4)
9:         Backpropagate: Compute gradients and update $\omega_{\ell,j}$ (model parameters $\theta$ remain frozen)

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable proficiency in in-context learning (ICL), where models adapt to new tasks through example-based prompts without requiring parameter updates. However, understanding how tasks are internally encoded and generalized remains a challenge. To address some of the empirical and technical gaps in the literature, we introduce an automated formulation for encoding task information in ICL prompts as a function of attention heads within the transformer architecture. This approach computes a single task vector as a weighted sum of attention heads, with the weights optimized causally via gradient descent. Our findings show that existing methods fail to generalize effectively to modalities beyond text. In response, we also design a benchmark to evaluate whether a task vector can preserve task fidelity in functional regression tasks. The proposed method successfully extracts task-specific information from in-context demonstrations and excels in both text and regression tasks, demonstrating its generalizability across modalities.
Learning Task Representations from In-Context Learning
Baturay Saglamthanks: Corresponding author: baturay.saglam@yale.edu, Xinyang Hu, Zhuoran Yang, Dionysis Kalogerias, Amin Karbasi Yale University
@@ 1 Introduction
Large language models (LLMs) based on the transformer architecture (Vaswani et al., 2017) have seen dramatic improvements in recent years. A notable feature of these models, such as GPT-3 (Brown et al., 2020), is their capability for in-context learning (ICL). This process involves the model receiving a prompt that includes demonstrations of a task in the form of input-output pairs. When presented with a new query input, the model can generate the appropriate output by extrapolating from the provided examples. For instance, after being prompted with a few examples, these models are capable of producing the antonyms of given input words. A concrete example is
[…]
@@
where the blue text is the prompt and the red text is the completion provided by the model. Not limited to linguistic tasks, transformers can also in-context learn a general class of functions $\mathcal{F}$ (Garg et al., 2023). Specifically, for any function $f\in\mathcal{F}$, the model is capable of approximating $f(x_{\text{query}})$ for a new query input $x_{\text{query}}$. This class may include linear or nonlinear relationships, which can be represented by various machine learning models such as linear regression, multi-layer ReLU networks, and decision trees.
@@
The ICL capability is particularly intriguing as it allows models to adapt to a wide range of downstream tasks on-the-fly—i.e., without requiring any parameter updates post-training (Brown et al., 2020). This indicates that models can extract abstract task information (the relationship between input-output pairs) from the prompt and use it to generate accurate responses for query inputs. However, due to the complex nature of the transformer architecture, the computational mechanisms that facilitate how models internally extract and encode task information are not well understood.
@@
Task vectors have been identified as key mechanisms for encoding task information during ICL (Todd et al., 2024; Li et al., 2024; Hendel et al., 2023; Liu et al., 2024). These vectors capture how models process prompts and embed the abstract in-context task as a numerical representation. While the effectiveness of these vectors is empirically well-supported, the precise computation of a task vector within the transformer architecture remains debated. Early findings suggest that layer or attention activations significantly influence ICL performance (Hendel et al., 2023; Liu et al., 2024; Todd et al., 2024).
However, there is no consensus on the optimal way to conceptualize or compute these vectors, resulting in varied methodologies that often excel in linguistics but fail to generalize to other tasks, such as functional regression. Since these methods rely on the internal components of transformers and architectural differences due to output modality (e.g., language head or numeric outputs) do not change their operation, any meaningful task representation should, in principle, be agnostic to output modality. Poor non-language performance therefore suggests that they may be capturing modality-specific patterns rather than a generalizable task representation. Additionally, these methods require adaptation to individual transformer architectures. For example, the layers over which the task vector is computed must be explicitly identified for each model (Li et al., 2024; Todd et al., 2024; Hendel et al., 2023).
@@
These limitations raise questions about the generalizability and practical utility of existing task vector formulations. To pursue a more standardized and automated pipeline for task encodings, we build on the recent finding that attention heads are the primary driver for vectorizing ICL tasks and hypothesize that:
@@ A New Benchmark for Regression Tasks
While prior work has explored where in-context tasks are represented, most studies focus solely on language, with limited attention to other modalities like functional regression. We introduce a new benchmark to assess task encodings in regression. This setup provides an intuitive framework for analyzing task encodings and ICL capabilities in regression by leveraging an out-of-distribution condition.
@@
Our method shows effective task encoding on the proposed regression benchmark and language datasets. Ablation studies indicate that this performance stems from aligning the distribution of the last hidden state with that of an optimally in-context-learned model, offering insights into how task steering is achieved.
For reproducibility, we provide the code in our GitHub repository [^1] https://github.com/baturaysaglam/ICL-task-repr.
@@ 2 Related Work
Initial studies on developing task representations for transformers are documented by Lampinen and McClelland (2020); Shao et al. (2023); Mu et al. (2023); Panigrahi et al. (2023); Ilharco et al. (2023). These works proposed methods to create compositional task encodings through model perturbations in the parameter space: “soft” prompts, codebooks, and meta-mappings. Notably, the term task vectors was first introduced by Ilharco et al. (2023). In contrast, through the application of causal mediation analysis (Pearl, 2001; Vig et al., 2020; Meng et al., 2022; Wang et al., 2023a; Geva et al., 2023), function vectors were discovered to exist inherently within the transformer architecture and to exhibit strong causal effects (Todd et al., 2024). This finding aligns with research on RNNs, where it was demonstrated that RNN hidden states can be grouped based on task similarities (Lake and Baroni, 2018; Hill et al., 2019).
Efforts to represent ICL tasks often derive a task vector from the layer activations associated with the dummy token following the input query (Hendel et al., 2023). This approach was later refined by averaging the layer activations of dummy tokens across a few-shot prompt and optimizing them (Li et al., 2024). However, these methods were primarily designed for language tasks, where input-output pairs are explicitly separated by specific tokens (e.g., $\rightarrow$”). This limitation restricts their applicability to broader domains, such as regression tasks. Another approach leverages the first principal component of the difference in layer activations to guide the ICL task, resulting in In-Context Vectors (ICVs) (Liu et al., 2024). Notably, attention heads have been argued to play a critical role in transferring information between token positions (Vaswani et al., 2017; Elhage et al., 2021b). Similarly, Todd et al. (2024) found that attention heads are primarily responsible for encoding ICL behavior. Their proposed formulation, Function Vector (FV), is computed as the sum of activations from a specific subset of attention heads, selected based on an indirect metric derived from causal inference literature (Pearl, 2001). In this study, we aim to develop a structured and automated method for extracting function vectors from transformer architectures, expanding on the approach of Todd et al. (2024).
@@ 3 Technical Preliminaries
We focus on decoder-only autoregressive transformer architectures.
3.1 Transformer Architecture
@@
$\operatorname{MultiHead}_{\ell}(Q_{\ell},K_{\ell},V_{\ell})=\operatorname{Concat}(\operatorname{head}_{\ell,1},\\ \ldots,\operatorname{head}_{\ell,J})W_{\ell}^{O},$
where $\mathbb{R}^{q}\ni\operatorname{head}_{\ell,j}\coloneqq\operatorname{Attn}_{\ell,j}=\operatorname{softmax}\big(Q_{\ell,j}K^{\top}_{\ell,j}/\sqrt{d_{k}})V_{\ell,j}$. Here $W_{\ell}^{O}\in\mathbb{R}^{Jq\times d}$ is the output projection matrix, and $Q_{\ell,j}$, $K_{\ell,j}$, and $V_{\ell,j}$ are the query, key, and value matrices for each attention head $j\leq J$ at layer $\ell$. The term $\sqrt{d_{k}}$ normalizes the softmax operation for stability, where $d_{k}$ is the dimension of the key matrix. This multi-head approach allows the model to dynamically adjust its focus across different parts of the input based on the context.
Note that each head at layer $\ell$ operates in a low-dimensional space $q<d$, distinct from the main hidden state residual stream of the transformer. As observed by Elhage et al. (2021a), the original transformer formulation (Vaswani et al., 2017) can be seen as dividing the matrix into a block form $W_{\ell,1},\ldots,W_{\ell,J}]$, followed by directly projecting each $\operatorname{head}{\ell,j}$ into the residual space. Consequently, the attention head output $a{\ell,j}$ can be defined using Vaswani et al.’s (2017) and Elhage et al.’s (2021a) notations as:
@@ 3.2 In-Context Learning
Regression tasks refer to a function class $\mathcal{F}$ that is learned through ICL. We closely follow the formulation proposed in (Garg et al., 2023) for training and testing models on regression data. For each prompt, a random function $f$ from $\mathcal{F}$ is sampled according to a distribution $\mathcal{D}_{\mathcal{F}}$, and a set of random inputs $x_{i}\in\mathbb{R}^{m}$ for $i=1,\ldots,T$ is drawn from $\mathcal{D}_{\mathcal{X}}$. These inputs are then evaluated by $f$ to produce the prompt $p^{f}=\{x_{1},f(x_{1}),\ldots,x_{T},f(x_{T}),x_{\text{query}}\}$. We denote the model’s output as $M_{\theta}(p^{f})$.
A function class $\mathcal{F}$ can be either linear or nonlinear. For linear functions, inputs are drawn from an isotropic Gaussian distribution $\mathcal{N}(0,I_{m})$, and a random function is defined by sampling a weight vector $w$ from $\mathcal{N}(0,I_{m})$, setting $f(x)=w^{\top}x$. Here, the task is characterized by the weight vector $w\in\mathbb{R}^{m}$. For nonlinear functions, possible forms of $f$ include multi-layer ReLU networks or decision trees. We employ the models pre-trained by Garg et al. (2023), with the training procedure and additional details provided in Appendix C.2.2.
@@
We focus on straightforward NLP applications such as antonym generation, with an example shown in (1). The model $M_{\theta}$ processes each prompt $p^{t}$ and produces a next-token distribution $M_{\theta}(\cdot\mid p^{t})$ over the vocabulary $\mathcal{V}$. In our evaluation of task representations, we follow Todd et al. (2024), who corrupted prompts $\tilde{p}_{i}^{t}$ by shuffling labels within each prompt. This shuffling disrupts the connection between inputs $x_{i,k}$ and outputs $\tilde{y}_{i,k}$, rendering the examples uninformative about the task and effectively “blocking” ICL.
Transformer Trans. +FV Trans. +ICV
(a) Linear
(b) Sparse linear
(c) 2-layer NN
@@ 4.1 Motivational Observations
Our study begins by noting that existing ICL task formulations are either inapplicable (Li et al., 2024; Hendel et al., 2023) (see Appendix C.2.6) or fail to generalize to regression tasks (Liu et al., 2024; Todd et al., 2024). To demonstrate this limitation, we design a new benchmark for functional regression.
The motivation for this benchmark is straightforward and follows standard practices from prior work on language tasks (see Section 3.2). We first prevent the model from extracting task understanding from in-context examples. If ICL performance is restored when the task vector is integrated into the model, it suggests that the task vector encodes the task information. However, techniques like shuffling labels or probing are unsuitable for regression tasks due to fundamental differences in how input-output relationships are represented and learned (see Appendices B.1 and B.2).
To suppress task-relevant information in regression prompts, we introduce an out-of-distribution (OOD) case. The model is initialized and pre-trained to handle sequences up to a maximum length $T_{\text{max}}$, ensuring it can process longer sequences. However, during pre-training, the model is exposed only to sequences of length up to $T_{\text{train}}<T_{\text{max}}$, leaving positional encodings for positions beyond $T_{\text{train}}$ untrained.
At test time, we evaluate the model on sequences longer than $T_{\text{train}}$ but within $T_{\text{max}}$. The focus is not on downstream accuracy but on robustness under OOD conditions. Since positional encodings beyond $T_{\text{train}}$ are untrained, the model cannot infer where the input and outputs are located within the prompt, and thus cannot extract the desired input-output relationship any further—the importance of positional information in ICL has also been considered by Wibisono and Wang (2024). However, if a representation contains task-relevant information (specific to the ideal function $f$), it should maintain functional task behavior and generalize to prompts longer than $T_{\text{train}}$.
The results, shown in Figure 1, reveal that the tested approaches fail to generalize across the evaluated class of functions, despite decent performance in the language domain. Because these methods rely on the internal components of transformers and are unaffected by architectural differences related to output modality (e.g., language heads or numeric outputs), any meaningful task representation should, in principle, remain consistent across output modalities.
As our investigation deepens, we articulate the following central research questions:
@@
    How can we develop a task representation in a principled manner that generalizes across modalities such as language and mathematical functions?
@@
    How can this representation be leveraged to preserve task fidelity on downstream tasks?
@@ 4.2 Learnable Task Vector
Our approach addresses key limitations of previous methods, such as Function Vectors, by automatically computing task formulations without the need for prior model analysis or manual adjustments.
@@
The FV encoding is based on strong causal evidence that attention heads are the main carriers of in-context task information (Todd et al., 2024). This formulation assumes equal contributions from all attention heads to the task representation by summing their activations. However, we argue that the influence of each head varies—some contribute significantly to task encoding, while others have a lesser impact. Constraining weights to unity is therefore impractical, and focusing exclusively on a subset of attention heads risks overlooking subtle contributions from others. To address this, attention heads should be weighted in the summation to accurately capture their varying contributions.
@@
Previous work typically integrates task encodings into models by adding vectors to the outputs of specific hidden layers through simple vector summation (Todd et al., 2024; Hendel et al., 2023; Li et al., 2024). However, this approach assumes uniform applicability across layers, ignoring that transformers’ hidden states represent progressively refined representations of the input. To overcome this limitation, we compute task vectors uniquely for each hidden layer, aligning them with the evolving representations. While initializing a unique set of attention head weights for each of the $L$ layers is an option, it would be computationally expensive. Instead, we compute layer-specific task vectors as a weighted sum of the attention heads within each layer:
@@
1:Input: Model $M_{\theta}$, task $t$, and number of samples NN
@@
3:Collect NN prompts $\{p^{t}_{i}\}_{i=1}^{N}$ demonstrating task $t$
@@
7:   Obtain the LTV-integrated model output: $\hat{y}_{\text{query},i}\leftarrow M_{\theta}(p^{t}_{i}\mid\mathbf{v}_{\Phi}^{t})$
8:   Compute the loss: $\mathcal{L}(y_{\text{query},i},\hat{y}_{\text{query},i})$ based on (3) or (4)
9:   Compute gradients and update $\omega_{\ell,j}$ (model parameters $\theta$ remain frozen)
@@
4.3 How to Learn Task Representations
We optimize LTV weights through gradient steps to approximate the “true” weights the model uses to represent ICL tasks, aiming to validate this formulation as a proof-of-concept. Consequently, this approach is not intended to directly address issues like length generalization shown in Figure 1. Instead, LTV serves as a mechanistic probe into the model’s internal structure, demonstrating that transformer architectures admit $d$-dimensional vectorization of task information that can be manipulated to steer behavior.
An effective strategy for learning the parameters $\Phi$ involves optimizing the LTV in scenarios where the model (with frozen parameters) fails to perform ICL, isolating the causal contribution of $\Phi$ to the task performance. This approach ensures our method is both data-driven and causal, as it directly attributes improvements to $|\Phi|=L\times J$ parameters, facilitating application to downstream tasks. The overall pipeline for training an LTV is summarized in Algorithm 1.
Regression Tasks
$\underset{\Phi}{\text{min}}\text{ }\mathbb{E}_{f\sim\mathcal{D}_{\mathcal{F}},x\sim\mathcal{D}_{\mathcal{X}}}\bigg[\Big(\tilde{M}_{\theta}\big(p^{f}\big.\mid\mathbf{v}_{\Phi}^{f}\big)-f(x_{\text{query}})\Big)^{2}\bigg],$ (3)
where $\mathbf{v}_{\Phi}^{f}=\{v_{\ell}^{f}\}_{\ell=1}^{L}$ represents the set of layer-wise LTVs computed for function $f$, $\tilde{M}_{\theta}(\cdot\mid\mathbf{v}_{\Phi}^{f})$ denotes the transformer output modified by integrating $\mathbf{v}_{\Phi}^{f}$ into the corresponding hidden layers, and the regression prompt $p^{f}$ is in the structure:
$p^{f}=\{x_{1},f(x_{1}),\ldots,x_{T_{\mathbf{v}}},f(x_{T_{\mathbf{v}}}),x_{\text{query}}\}.$
Language Tasks
$\underset{\Phi}{\text{min}}\text{ }\mathbb{E}_{\tilde{p}^{t}\sim\tilde{P}^{t}}\Big[-\log\Big(\tilde{M}_{\theta}\big(y_{\text{query}}\mid\tilde{p}^{t};\mathbf{v}_{\Phi}^{t}\big)\Big)\Big],$ (4)